OpenStudy (anonymous):

PLEASE HELP Use the rational zero theorem to find all the real zero of the polynomial function. Use the zeros to factor f over the real numbers. f(x)=2x^4+x^3-15x^2-7x+7 x=_____ OR There is no real zero Use the real zeros to factor f f(x)=________ (type answer in factored form)

OpenStudy (anonymous):

Do you know how to use the rational root test?

OpenStudy (anonymous):

no. not that I remember

OpenStudy (anonymous):

Basically what it is is that if a polynomial has rational roots, they can be written in the form p/q where p is an integer factor of the constant term, and q is an integer factor of the leading coefficient So in your case, if it has any rational roots, they will be one of \[\pm \frac{1,7}{1,2}\] which is basically shorthand for \[-1, 1, -7, 7, -\frac{1}{2}, \frac{1}{2}, -\frac{7}{2}, \frac{7}{2}\] So these are your possible roots, all you have to do is plug them in, and if you get zero it is a root :)

OpenStudy (anonymous):

Then once you have found a root \(r\), you can use synthetic division to factor it out by dividing the polynomial by \(x - r\)

OpenStudy (anonymous):

okay?

OpenStudy (anonymous):

How you found any zero's in the above list?

OpenStudy (anonymous):

still confused

OpenStudy (anonymous):

Basically the rational root theorem tells us that if there are any rational roots, they have to be one of these: \[-1, 1, -7, 7, -\frac{1}{2}, \frac{1}{2}, -\frac{7}{2}, \frac{7}{2}\] You can check if these are roots by plugging them into the function and finding out if they equal zero, for example \[\text{f}(x)=2x^4+x^3-15x^2-7x+7\] \[\text{f}(-1) = 2(-1)^4 + (-1)^3 - 15(-1)^2 - 7(-1) + 7 =\] \[2 - 1 - 15 + 7 + 7 = 0\] So since f(-1) = 0, -1 is a root of the polynomial There is another rational root in that list, x = \(1 \over 2\), and the others aren't rational roots So we can use syntetic division to factor out \((x + 1)\) and \((x - {1 \over 2})\)

OpenStudy (anonymous):

x= -1 and (1/2)

OpenStudy (anonymous):

Yes, so use long division / synthetic division to factor them both out, then check if there are any other real zero's by solving the quadratic factor that's left

OpenStudy (anonymous):

how would you start with the synthetic division?

OpenStudy (anonymous):

It's a bit hard to type it here, and I'm terrible at drawing so the best way to explain it is just give you a link: http://www.youtube.com/watch?v=bZoMz1Cy1T4 If you don't understand you can ask here :)

OpenStudy (anonymous):

so I know how to start it now but when I did it I got -4 as a remainder for using -1

OpenStudy (anonymous):

ok let me draw it  There should be no remainder \[2x^4 + x^3 - 15x^2 - 7x + 7 = (x + 1)(2x^3 - x^2 -14x + 7)\]

OpenStudy (anonymous):

opps did my wrong! then I use the quadratic factor?? and those will be my x? how will I find the real zeros??

OpenStudy (anonymous):

for the zero \(\frac{1}{2}\) you can repeat the process to get \[(x + 1)(x - {1 \over 2})(2x^2 - 14)=(x + 1)(2x - 1)(x^2 - 7)\] The other two roots are if \(x^2 - 7 = 0\), and we already know that these are irrational because if they were rational we would've found them using the rational root theorem :)

OpenStudy (anonymous):

how will we find the real zero of the factor of f?

OpenStudy (anonymous):

set it equal to zero: \[\begin{eqnarray} (x + 1) = 0 & \text{gives} & x = -1\\ (2x - 1) = 0 & \text{gives} & x = {1 \over 2}\\ (x^2 - 7) = 0 & \text{gives} & x = \pm\sqrt{7} \end{eqnarray}\] So the four real roots are -1, \(1 \over 2\), \(\sqrt 7\) and \(-\sqrt 7\).

OpenStudy (anonymous):

okay so then it would be the same with the x on the first question.

OpenStudy (anonymous):

yes, the four real zeros are x = -1, x = 1/2, x = sqrt(7) and x = -sqrt(7)

OpenStudy (anonymous):

so basically its the same thing.

OpenStudy (anonymous):

yeah, a zero is the same as a root

OpenStudy (anonymous):

so I need my answer in factored from or radicals as needed. would squ. root of 7 still count. sorry asking a lot of questions

OpenStudy (anonymous):

I would leave the factored form as (x + 1)(2x - 1)(x^2 - 7), because radicals don't look that pretty in factored form. They might want you to also factor (x^2 - 7) as (x - sqrt(7))(x + sqrt(7)), but I wouldn't do it like that myself.

OpenStudy (anonymous):

okay thank you!!!!!!